OpenStudy (anonymous):

If f(x) = 0.8(2 – x), what is the value of f(–2)?

OpenStudy (futureescape):

3.2, I think

OpenStudy (jadeishere):

distribute -2 where 'x' variable are; f(x) = 0.8(2 - (-2)) f(x) = 0.8(2 + 2) f(x) = 0.8(4) f(x) = 3.2

OpenStudy (muzzack):

@Jadeishere is right but in explanation time: when you are replacing a variable in a fucntion, it changes all the variables into the desired number so in this case you would replace x=-2 then you do your eqaution now what if i say i want a sequence... well thats different

OpenStudy (anonymous):

Can anyone here help me?

OpenStudy (anonymous):

Thank you all

OpenStudy (jadeishere):

Yeah ^^^^

OpenStudy (jadeishere):

I can try >.< different question though, not on this one